# Limit Computation Over Posets via Minimal Initial Functors

### v1

@@ 3.4. Computing Limits via Initial Scaffolds
Let $Q$ be either a finite poset or an interval in $\mathbb{N}^{d}$, and let $G\colon Q\to\mathbf{Vec}$ be a functor. In this paper, computing $\lim G$ means computing a presection basis of $G$ (Definition˜2.7).
@@
    While this definition of limit computation does not entail explicit computation of the cone maps of $\lim G$, Proposition˜2.6 makes clear that the cone maps are readily obtained from what we compute: Given a presection basis of $\lim G$, each cone map to a minimum is given by coordinate projection, while each cone map to a non-minimum is given by composing a coordinate projection with a structure map of $G$.
@@
    It is easily checked that if $j\colon P\hookrightarrow Q$ is an initial scaffold of $Q$, then the presection bases of $\lim G$ and $\lim(G\circ j)$ are identical. Therefore, according to Definition˜3.18, the problems of computing $\lim G$ and $\lim(G\circ j)$ are identical.
@@
    a choice of ordered basis for each of the vector spaces $(G_{q})_{q\in Q}$ and
@@
    the matrices $G_{pq}]_{p\leq q\in Q}$ representing the structure maps $(G_{pq})_{p\leq q\in Q}$ with respect to these bases.
@@
Proposition 3.21 ([jeannerod2013rank]).
A $y\times z$ matrix $A$ of rank $r$ can be transformed to row echelon in time $O(yzr^{\omega-2})$. Hence, using the standard $O(z^{2})$ backsolve algorithm, a basis for $\ker(A)$ can be computed in time $O(yz^{\omega-1})$.
We refer to the algorithm underlying Proposition˜3.21 as Gaussian Elimination, though it is different than the usual cubic-time version of Gaussian Elimination.
Together, Proposition˜3.21 and Proposition˜2.6 yield the following naive bound on the cost of limit computation, where
@@
The solution to the system of equations in Proposition˜2.6 is the kernel of a linear system with $O(r|Q|)$ equations and $O(rn)$ variables. The result now follows from Proposition˜3.21. ∎
We can improve on Proposition˜3.22 by computing initial scaffolds. Specifically, given a poset $Q$ as in either Theorem˜3.13 or Theorem˜3.17 and $Q$-module $G$, we compute $\lim G$ in three steps:
@@
    Compute an initial scaffold $j\colon P\hookrightarrow Q$ using the algorithm of Theorem˜3.13 or Theorem˜3.17.
@@
    Compute a matrix representation of $G\circ j$.
@@
    Compute $\lim(G\circ j)$ by applying Proposition˜3.22.
Let $\Gamma$ denote the cost of the second step.
@@
Given the Hasse diagram $(V,E)$ of a finite poset $Q$ with $n$ minima and a $Q$-module $G$ with $r(G)=r$, the above approach computes $\lim G$ in time
$O\left(|V||E|+|P|n^{\omega-1}r^{\omega}\right)+\Gamma.$
@@
By Theorem˜3.13, we can compute an initial scaffold of $Q$ in time $O(|V||E|)$. The result now follows from Proposition˜3.22. ∎
@@
Given the upset presentation of an interval $Q\subseteq\mathbb{N}^{d}$ with $\|Q\|=s$ and a $Q$-module $G$ with $r(G)=r$, the above approach computes $\lim G$ in time
@@
    $O(r^{\omega}s^{\omega+1})+\Gamma$ for $d>3$.
@@
By Theorem˜3.17, we can compute an initial scaffold $j\colon P\hookrightarrow Q$ in time $O(s\log s)$ for $d=2,3$ and $O(s^{4})$ for $d>3$. In addition, Theorem˜3.10 tells us that $|P|=\Theta(s)$ if $d=2,3$ and $|P|=\Theta(s^{2})$ if $d>3$. The result now follows from Proposition˜3.22. ∎
The cost $\Gamma$ depends on the format in which the input diagram $G$ is given. For example, if we are given a matrix representation of $G$, then $\Gamma$ is negligible. However, in the TDA applications that initially motivated this work, we typically are not given matrix representation of $G$; instead, we are given a free presentation of $G$ or, more generally, a chain complex of free functors whose homology is isomorphic to $G$. In this case, we may apply Proposition˜3.26 below.
@@
We can compute a matrix representation of $H\circ i$ in time $O(|P|r^{\omega})$.
We prove Proposition˜3.26 in Section˜7.1 via straightforward linear algebra.
@@
Note that if $g=0$, then the chain complex of Proposition˜3.26 is in fact a free presentation of $H$. In the general case where $g$ may be non-zero, one can in principle first compute a (minimal) presentation of $H$, and then take this as the input to algorithm underlying Proposition˜3.26. In TDA applications, it has been observed that the size of a minimal presentation of $H$ is often far smaller than the size of the original chain complex [fugacci2023compression].
In the case $Q=\mathbb{N}^{2}$ and in the case of $0^{\mathrm{th}}$ simplicial homology of filtrations indexed by arbitrary posets, efficient algorithms are known for computing a minimal presentation of $H$ [morozov2025computing, lesnick2022minimal, fugacci2023compression, bauer2023efficient]. In more general settings, the problem can be solved via Gröbner basis algorithms, e.g., Schreyer’s algorithm and its variants [schreyer1980berechnung, cox1998using, erocal2016refined, la1998strategies], but existing open-source implementations of such approaches are expensive on the types of large, sparse input considered in TDA [lesnick2022minimal].
@@
While some TDA constructions yield a chain complex whose chain modules are not free [botnan2023introduction, Section 5], such a chain complex can often readily be converted to a free chain complex with isomorphic homology [lesnick2015interactive, chacholski2017combinatorial].
@@
Given the Hasse diagram $(V,E)$ of a finite poset $Q$, we can compute a $|V|\times|V|$ binary matrix explicitly representing the partial order on $Q$ in time $O(|V||E|)$, by performing a breadth-first search starting from each $v\in V$. Using this matrix, we can check whether any relation $p\leq q$ holds in constant time, as required in Proposition˜3.26.
Together, Corollaries˜3.23, 3.13, 3.29 and 3.26 imply the following:
@@
Similarly, Corollaries˜3.24, 3.17 and 3.26 imply items (ii) and (iii) of the following corollary, which is an analogue of Corollary˜3.30 for intervals in $\mathbb{N}^{d}$.
@@
Given an $(\mathbb{N}^{d},r)$-complex with homology $H$ and the upset presentation of an interval $i\colon Q\hookrightarrow\mathbb{N}^{d}$ with $s=\|Q\|$, we can compute $\lim(H\circ i)$} in time
@@
    $O(r^{\omega}s^{\omega+1})$ for $d>3$.
We prove Corollary˜3.31 (i) in Section˜7.2 using a somewhat different approach than for parts (ii) and (iii): By Example˜3.4, an interval in $\mathbb{N}^{2}$ has a unique initial scaffold, which is a zigzag poset. This allows us to compute the limit using zigzag persistence computation, rather than the general limit computation of Proposition˜3.22, and this turns out to be more efficient.
@@
Corollaries˜3.31 and 3.24 also hold if $Q$ is instead specified by its minima and maxima and $s$ is the total number of these.
@@ 3.5. Generalized Rank Computation
As an application of our main results, we consider the problem of computing the generalized rank of a diagram of vector spaces, a problem arising in TDA which has previously been studied in [DKM22, dey2024computing, asashiba2024interval].
Definition 3.34 ([kim2018generalized]).
@@
    The invariant $\operatorname{grank}(G)$ has a simple representation-theoretic interpretation: It is the number of interval summands of $G$ with support $Q$ [chambers2018persistent, Corollary 3.2].
@@
    Recent work by Xin et al. [xin2023gril] has used the generalized ranks along certain subintervals of $\mathbb{N}^{2}$ called worms to define GRIL, a novel vectorization of 2-parameter persistence for supervised learning.
@@
    Applying Möbius inversion to the function sending an interval $j\colon Z\hookrightarrow Q$ to $\operatorname{grank}(G\circ j)$ yields an invariant of $G$ called the generalized persistence diagram [kim2018generalized], a generalization of the usual persistence diagram of a $\mathbb{N}$-module that has been actively studied in recent work [botnan2024signed, kim2024bigraded, kim2025super, carriere2025sparsification, asashiba2023approximation]. The large size of generalized persistence diagrams is an obstacle to their practical use [kim2025super], but they are of interest from a theoretical standpoint.
Before stating our results on generalized rank computation, we briefly discuss prior work on this. Dey, Kim, and Mémoli [DKM22] gave the following bound on the cost of computing a generalized rank over an interval in $\mathbb{N}^{2}$:
Theorem 3.35 ([DKM22]).
Given an $(\mathbb{N}^{2},r)$-complex with homology $H$ and the $s$ extrema of a finite interval $i\colon Q\hookrightarrow\mathbb{N}^{2}$, we can compute $\operatorname{grank}(H\circ i)$ in time $O((s+r)^{\omega})$.
The algorithm underlying Theorem˜3.35 determines $\operatorname{grank}(G)$ by computing the zigzag barcode of $G\circ j$, where $j\colon P\hookrightarrow Q$ is a zigzag containing the initial and final scaffolds of $Q$. It is shown in [DKM22] that $\operatorname{grank}(G)$ is the number of copies of the full interval $P$ in this barcode.
Inspired by Theorem˜3.35, Dey and Xin [dey2024computing] gave an algorithm to compute the generalized rank over a general poset by reducing the problem to a zigzag persistence computation. However, it is actually asymptotically more efficient to compute the generalized rank by directly computing the limits and colimits, even without the speedups enabled by initial and final functors.
For $Q$ a finite poset, $i\colon P\hookrightarrow Q$ an interval, and $G$ a $Q$-module, Asashiba and Liu [asashiba2024interval] gave an explicit formula for $\operatorname{grank}(G\circ i)$ as the difference of the rank of two matrices. In that work, the formalism of (co)limits is not explicitly used, and a bound on the cost of computing $\operatorname{grank}(G\circ i)$ is not explicitly given.
@@
Let $Q$ be a finite, connected poset with $s$ extrema and let $^{\mathcal{I}}, PP^{\mathcal{F}}$ denote initial and final scaffolds of $Q$. Given the Hasse diagram $(V,E)$ of $Q$ and a $(Q,r)$-complex with homology $H$, we can compute $\operatorname{grank}(H)$ in time
$O(|V||E|+(|P^{\mathcal{I}}|+|P^{\mathcal{F}}|)s^{\omega-1}r^{\omega}).$
@@
Given an $(\mathbb{N}^{d},r)$-complex with homology $H$ and the $s$ extrema of a finite interval $i\colon Q\hookrightarrow\mathbb{N}^{d}$, we can compute $\operatorname{grank}(H\circ i)$ in time
@@
    $O(r^{\omega}s^{\omega+1})$ for $d>3$.
Note that Corollary˜3.37 (i) slightly strengthens Theorem˜3.35. As we explain in Section˜8, Corollary˜3.36 and Corollary˜3.37 (ii) and (iii) follow readily from our main results, via straightforward linear algebra. To prove Corollary˜3.37 (i), we use an additional argument to cast the generalized rank computation as a zigzag persistence computation. This argument is similar to the one in [DKM22], though we frame it in the general formalism of initial and final scaffolds.
@@ 4. Proof of Structure Theorem for Minimal Initial Functors
and the empty set is disconnected (see Definition˜2.2), we have $M_{Q}\subseteq I_{Q}$. Let $E_{Q}=I_{Q}\setminus M_{Q}$, i.e.,
@@
We note that $\operatorname{im}F$ is not necessarily a subcategory of $D$, since $F(\hom C)$ may not be closed under composition; see Example˜4.4 below. We sometimes abuse notation slightly and, e.g., write $d\in\operatorname{im}F$ to mean $d\in F(\operatorname{Ob}C)$, or $\gamma\in\operatorname{im}F$ to mean $\gamma\in F(\hom C)$.
Since the definition of a connected category (Definition˜2.2) does not involve composition of morphisms, it extends to any collections of objects and morphisms in an ambient category. In particular, the notion of connectivity makes sense for $\operatorname{im}F$. We use this to state the next proposition, whose straightforward proof we omit.
@@
If a category $C$ is connected and $F\colon C\to D$ is any functor, then $\operatorname{im}(F)$ is also connected.
We let $\overline{\operatorname{im}F}$ denote the subcategory of $D$ generated by $\operatorname{im}F$, i.e., the category obtained from $\operatorname{im}F$ by including the compositions of all finite chains of composable morphisms in $\operatorname{im}F$.
@@
To prove (i), first recall from Section˜2.3 that for any functor $F\colon C\to Q$ and $q\in Q$, $(F\downarrow q)$ can be identified with the full subcategory of $C$ with objects $\{c\in\operatorname{Ob}c\mid F(c)\leq q\}$. If $F$ is initial, then by definition $(F\downarrow q)$ is connected, hence non-empty, so $M_{Q}\subseteq\operatorname{im}(F)$.
Next we show that $E_{Q}\subseteq\operatorname{im}F$. To do so, we assume to the contrary that there exists $q\in E_{Q}$ with $q\not\in\operatorname{im}F$, and show that then $(F\downarrow q)$ is disconnected, a contradiction. Let $\hat{F}\colon(F\downarrow q)\to Q$} be the restriction of $F$ to $(F\downarrow q)$ and note that $\operatorname{im}\hat{F}=\operatorname{im}F\cap\operatorname{dn}[q]$. Since we assume $q\not\in\operatorname{im}F$, we have $\operatorname{im}F\cap\operatorname{dn}[q]\subseteq\operatorname{dn}(q)$. Moreover, $\operatorname{im}F$ contains all minima of $Q$, so since $Q$ has finite downsets, $\operatorname{im}F$ has non-empty intersection with each component of $\operatorname{dn}(q)$. Since $q$ is essential, $\operatorname{dn}(q)$ is disconnected. It follows that $\operatorname{im}\hat{F}=\operatorname{im}F\cap\operatorname{dn}(q)$ is disconnected. Thus, $(F\downarrow q)$ is disconnected by the contrapositive of Proposition˜4.2, a contradiction. We conclude that $E_{Q}\subseteq\operatorname{im}F$, and hence that $I_{Q}\subseteq F(\operatorname{Ob}C)$, establishing (i).
In view of (i), proving (ii) amounts to showing that for each $q\in E_{Q}$ and component $A$ of $\operatorname{dn}(q)$, there is a relation $m<q$ in $\overline{\operatorname{im}F}$ with $m$ a minimum in $A$. Since $Q$ has finite downsets and $M_{Q}\subseteq I_{Q}\subseteq\operatorname{im}F$ by (i), we have $\operatorname{im}F\cap A\neq\emptyset$. Moreover, since $(F\downarrow q)$ is connected, Proposition˜4.2 implies that $\operatorname{im}\hat{F}=\operatorname{im}F\cap\operatorname{dn}[q]$ is connected. Thus, since $q\in\operatorname{im}F$, there exists at least one relation $a_{0}<q$ in $\operatorname{im}F$ where $a_{0}\in A$. If $a_{0}\in M_{Q}$, we are done. Otherwise, by applying the same argument to any component of $\operatorname{dn}(a_{0})$, we obtain a relation $a_{1}<a_{0}$ in $\operatorname{im}F$ with $a_{1}\in A$; indeed to show the existence of $a_{0}$, we did not use the fact that $q$ is essential, so the same argument applies. Since $Q$ has finite downsets, by continuing in this way, we eventually obtain a string of relations
@@
The following example shows that Proposition˜4.3 (ii) is no longer true if we replace $\overline{\operatorname{im}F}$ with $\operatorname{im}F$ in the statement.
@@
Figure 2. The initial functor $F\colon R\to Q$ of Example˜4.4 maps each element of $R$ to the element of $Q$ directly below it. For example, $F(b)=F(c)=x$.
@@
Given any category $C$, there is an associated thin category $\operatorname{thin}(C)$ with the same objects such that $\hom(c,c^{\prime})$ is empty in $\operatorname{thin}(C)$ if and only if $\hom(c,c^{\prime})$ is empty in $C$. Moreover, we have a unique functor $\pi\colon C\to\operatorname{thin}(C)$ which is the identity on objects.
If $D$ is a thin category (say a poset), then any functor $F\colon C\to D$ induces a functor $\operatorname{thin}(F)\colon\operatorname{thin}(C)\to D$ satisfying $F=\operatorname{thin}(F)\circ\pi$.

### v2

@@ 3.4. Computing Limits via Initial Scaffolds
Let $Q$ be either a finite poset or an interval in $\mathbb{N}^{d}$, and let $G\colon Q\to\mathbf{Vec}$ be a functor. In this paper, computing $\lim G$ means computing a presection basis of $G$ (Definition˜2.11).
@@
    While this definition of limit computation does not entail explicit computation of the cone maps of $\lim G$, Proposition˜2.10 makes clear that the cone maps are readily obtained from what we compute: Given a presection basis of $\lim G$, each cone map to a minimum is given by coordinate projection, while each cone map to a non-minimum is given by composing a coordinate projection with a structure map of $G$.
@@
    Using Proposition˜2.14, it is easily checked that if $j\colon P\hookrightarrow Q$ is an initial scaffold of $Q$, then the presection bases of $\lim G$ and $\lim(G\circ j)$ are identical. Therefore, according to Definition˜3.17, the problems of computing $\lim G$ and $\lim(G\circ j)$ are identical.
@@
    a choice of ordered basis for each of the vector spaces $(G_{q})_{q\in Q}$,
@@
    the matrix representation $G_{pq}]$ of each structure map $G_{pq}$ of $G$, with respect to these bases.
@@
Proposition 3.20 ([25]).
A $y\times z$ matrix $A$ of rank $r$ can be transformed to row echelon in time $O(yzr^{\omega-2})$. Hence, using the standard $O(z^{2})$-time backsolve algorithm, a basis of $\ker(A)$ can be computed in time
$O(yzr^{\omega-2}+z^{2})=O(yz^{\omega-1}).$
We refer to the algorithm underlying Proposition˜3.20 as Gaussian Elimination, though it is different than the usual cubic-time version of Gaussian Elimination.
Together, Proposition˜3.20 and Proposition˜2.10 yield the following naive bound on the cost of limit computation, where
@@
The solution to the system of equations in Proposition˜2.10 is the kernel of a linear system with $O(r|Q|)$ equations and $O(rn)$ variables. The result now follows from Proposition˜3.20. ∎
We aim to refine Proposition˜3.21 by computing initial scaffolds. To this end, given a poset $Q$ which is either finite or an interval in $\mathbb{N}^{d}$, as well as a $Q$-module $G$, we compute $\lim G$ in three steps:
@@
    compute an initial scaffold $j\colon P\hookrightarrow Q$ via Theorem˜3.13 or Theorem˜3.16.
@@
    compute a matrix representation of $G\circ j$,
@@
    compute $\lim(G\circ j)$ by applying Proposition˜3.21.
Letting $\Gamma$ denote the cost of the second step, this approach leads to the following bounds on the cost of limit computation:
@@
Given the Hasse diagram $(V,E)$ of a finite poset $Q$ with $n$ minima and a $Q$-module $G$ with $r(G)=r$, we can compute $\lim G$ in time
$O\left(|V||E|+|P|n^{\omega-1}r^{\omega}\right)+\Gamma,$
where $P$ is an initial scaffold of $Q$.
@@
By Theorem˜3.13, we can compute an initial scaffold $P\subseteq Q$ in time $O(|V||E|)$. The result now follows from Proposition˜3.21. ∎
@@
Given the upset presentation of an interval $Q\subseteq\mathbb{N}^{d}$ with $\|Q\|=n$ and a $Q$-module $G$ with $r(G)=r$, we can compute $\lim G$ in time
@@
    $O(n^{4}+n^{\omega+1}r^{\omega})+\Gamma$ for $d>3$.
@@
By Theorem˜3.16, we can compute an initial scaffold $P\subseteq Q$ in time $O(n\log n)$ for $d=2,3$ and $O(n^{4})$ for $d>3$. In addition, Theorem˜3.10 tells us that $|P|=\Theta(n)$ if $d=2,3$ and $|P|=\Theta(n^{2})$ if $d>3$. The result now follows from Proposition˜3.21. ∎
The cost $\Gamma$ depends on the format in which the input diagram $G$ is given. For example, if we are given a matrix representation of $G$, then $\Gamma$ is negligible. However, in the TDA applications that initially motivated this work, we typically are not given a matrix representation of $G$; instead, we are given a free presentation of $G$ or, more generally, a chain complex of free functors whose homology is isomorphic to $G$. In this case, we may apply Proposition˜3.25 below.
@@
Then we can compute a matrix representation of $H\circ i$ in time $O(|P|r^{\omega})$.
We prove Proposition˜3.25 in Section˜7.1 via straightforward linear algebra.
@@
Note that if $g=0$, then the $(Q,r)$-complex of Proposition˜3.25 is in fact a free presentation of $H$. In the general case where $g$ may be non-zero, one can in principle first compute a (minimal) presentation of $H$, and then take this as the input to algorithm underlying Proposition˜3.25.
In the case $Q=\mathbb{N}^{2}$ and in the case of $0^{\mathrm{th}}$ simplicial homology of filtrations indexed by arbitrary posets, efficient algorithms are known for computing a minimal presentation of $H$ [37, 31, 23, 4]. In the case $Q=\mathbb{N}^{2}$, it has been observed that the size of a minimal presentation of $H$ is often far smaller than the size of the original chain complex [23]. For $Q=\mathbb{N}^{d}$, a minimal presentation of $H$ can be computed via Gröbner basis algorithms, e.g., Schreyer’s algorithm and its variants [40, 15, 22, 29], but existing open-source implementations of such algorithms are expensive on the types of large, sparse input considered in TDA [31].
@@
While some TDA constructions yield a chain complex whose chain modules are not free [7, Section 5], such a chain complex can often readily be converted to a free chain complex with isomorphic homology [30, 12, 3].
@@
Given the Hasse diagram $(V,E)$ of a finite poset $Q$, in time $O(|V||E|)$ we can compute a $|V|\times|V|$ binary matrix that explicitly represents the partial order on $Q$, by performing a breadth-first search starting from each $v\in V$. Using this matrix, we can check whether any relation $p\leq q$ holds in constant time, as required in Proposition˜3.25.
Together, Corollaries˜3.22, 3.13, 3.28 and 3.25 imply the following:
@@
Similarly, Corollaries˜3.23, 3.16 and 3.25 imply items (ii) and (iii) of the following corollary, which is an analogue of Corollary˜3.29 for intervals in $\mathbb{N}^{d}$.
@@
Given an $(\mathbb{N}^{d},r)$-complex with homology $H$ and the upset presentation of an interval $i\colon Q\hookrightarrow\mathbb{N}^{d}$ with $\|Q\|=n$, we can compute $\lim(H\circ i)$} in time
@@
    $O(n^{4}+n^{\omega+1}r^{\omega})$ for $d>3$.
We prove Corollary˜3.30 (i) in Section˜7.2 using a somewhat different approach than for parts (ii) and (iii): By Example˜3.4, an interval in $\mathbb{N}^{2}$ has a unique initial scaffold, which is a zigzag poset. This allows us to compute the limit using a zigzag persistence computation rather than the general limit computation of Proposition˜3.21, and this turns out to be more efficient.
@@
Corollaries˜3.30 and 3.23 both also hold if $Q$ is instead specified by its minima and maxima and $n$ is the total number of these.
@@ 3.5. Generalized Rank Computation
As an application of our main results, we consider the problem of computing the generalized rank of a diagram of vector spaces, a problem arising in TDA that has previously been studied in [16, 17, 2].
Definition 3.33 ([27]).
@@
    The invariant $\operatorname{grank}(G)$ has a simple representation-theoretic interpretation: It is the number of interval summands of $G$ with support $Q$ [13, Corollary 3.2].
@@
    Recent work by Xin et al. [41] has used the generalized ranks along certain subintervals of $\mathbb{N}^{2}$ called worms to define GRIL, a novel vectorization of 2-parameter persistence for supervised learning.
@@
    Applying Möbius inversion to the function sending an interval $j\colon Z\hookrightarrow Q$ to $\operatorname{grank}(G\circ j)$ yields an invariant of $G$ called the generalized persistence diagram [27], a generalization of the usual persistence diagram of a $\mathbb{N}$-module, which has been actively studied in recent work [6, 28, 26, 11, 1]. The large size of generalized persistence diagrams is an obstacle to their practical use [26], but they are of interest from a theoretical standpoint.
Before stating our results on generalized rank computation, we briefly discuss prior work on this. Dey, Kim, and Mémoli [16] gave the following bound on the cost of computing a generalized rank over an interval in $\mathbb{N}^{2}$:
Theorem 3.34 ([16]).
Given an $(\mathbb{N}^{2},r)$-complex with homology $H$ and the $n$ extrema of a finite interval $i\colon Q\hookrightarrow\mathbb{N}^{2}$, we can compute $\operatorname{grank}(H\circ i)$ in time $O((n+r)^{\omega})$.
The algorithm underlying Theorem˜3.34 determines $\operatorname{grank}(G)$ by computing the zigzag barcode of $G\circ j$, where $j\colon P\hookrightarrow Q$ is a zigzag containing the initial and final scaffolds of $Q$. It is shown in [16] that $\operatorname{grank}(G)$ is the number of copies of the full interval $P$ in this barcode.
Inspired by Theorem˜3.34, Dey and Xin [17] gave an algorithm to compute the generalized rank over a general poset by reducing the problem to a zigzag persistence computation. However, it turns out to be asymptotically more efficient to compute the generalized rank by directly computing the limits and colimits; this is true even if one does not exploit the speedups enabled by initial and final functors.
For $Q$ a finite poset, $i\colon P\hookrightarrow Q$ an interval, and $G$ a $Q$-module, Asashiba and Liu [2] gave an explicit formula for $\operatorname{grank}(G\circ i)$ as the difference of the rank of two matrices. In that work, the formalism of (co)limits is not explicitly used, and a bound on the cost of computing $\operatorname{grank}(G\circ i)$ is not explicitly given.
@@
Let $Q$ be a finite, connected poset with $n$ extrema and let $^{\mathcal{I}}, PP^{\mathcal{F}}$ be initial and final scaffolds of $Q$. Given the Hasse diagram $(V,E)$ of $Q$ and a $(Q,r)$-complex with homology $H$, we can compute $\operatorname{grank}(H)$ in time
$O(|V||E|+(|P^{\mathcal{I}}|+|P^{\mathcal{F}}|)n^{\omega-1}r^{\omega}).$
@@
Given an $(\mathbb{N}^{d},r)$-complex with homology $H$ and the $n$ extrema of a finite interval $i\colon Q\hookrightarrow\mathbb{N}^{d}$, we can compute $\operatorname{grank}(H\circ i)$ in time
@@
    $O(n^{4}+n^{\omega+1}r^{\omega})$ for $d>3$.
Note that Corollary˜3.36 (i) slightly strengthens Theorem˜3.34. As we explain in Section˜8, Corollary˜3.35 and Corollary˜3.36 (ii) and (iii) follow readily from our main results, via straightforward linear algebra. To prove Corollary˜3.36 (i), we use an additional argument to cast the generalized rank computation as a zigzag persistence computation. This argument is similar to the one in [16], though we frame it in the general formalism of initial and final scaffolds.
@@ 4. Proof of Structure Theorem for Minimal Initial Functors
and the empty set is disconnected (see Definition˜2.3), we have $M_{Q}\subseteq I_{Q}$. Let $E_{Q}=I_{Q}\setminus M_{Q}$, i.e.,
@@
We sometimes abuse notation slightly and, e.g., write $d\in\operatorname{im}F$ to mean $d\in F(\operatorname{Ob}C)$, or $\gamma\in\operatorname{im}F$ to mean $\gamma\in F(\hom C)$.
We note that $\operatorname{im}F$ is not necessarily a subcategory of $D$, since $F(\hom C)$ may not be closed under composition; see Example˜4.5 below. We let $\overline{\operatorname{im}F}$ denote the subcategory of $D$ generated by $\operatorname{im}F$, i.e., the category obtained from $\operatorname{im}F$ by including the compositions of all finite chains of composable morphisms in $\operatorname{im}F$.
We omit the straightforward proof of the following result:
Lemma 4.2.
If $F$ is an embedding (see Definition˜2.1), then $\operatorname{im}F$ is closed under composition of morphisms, i.e., $\overline{\operatorname{im}F}=\operatorname{im}F$.
Since the definition of a connected category (Definition˜2.3) does not involve composition of morphisms, it extends to any collections of objects and morphisms in an ambient category. In particular, the notion of connectivity makes sense for $\operatorname{im}F$. We use this to state the next proposition, whose straightforward proof we omit.
@@
If a category $C$ is connected and $F\colon C\to D$ is any functor, then $\operatorname{im}F$ is also connected.
@@
To prove (i), first recall from Section˜2.3 that for any functor $F\colon C\to Q$ and $q\in Q$, $(F\downarrow q)$ can be identified with the full subcategory of $C$ with objects $\{c\in\operatorname{Ob}C\mid F(c)\leq q\}$. If $F$ is initial, then by definition $(F\downarrow q)$ is connected, hence non-empty, so $M_{Q}\subseteq\operatorname{im}F$.
Next we show that $E_{Q}\subseteq\operatorname{im}F$. To do so, we assume to the contrary that there exists $q\in E_{Q}$ with $q\not\in\operatorname{im}F$, and show that then $(F\downarrow q)$ is disconnected, a contradiction. Let $\hat{F}\colon(F\downarrow q)\to Q$} be the restriction of $F$ to $(F\downarrow q)$ and note that $\operatorname{im}\hat{F}=\operatorname{im}F\cap\operatorname{dn}[q]$. Since we assume $q\not\in\operatorname{im}F$, we have $\operatorname{im}F\cap\operatorname{dn}[q]\subseteq\operatorname{dn}(q)$. Moreover, $\operatorname{im}F$ contains all minima of $Q$, so since $Q$ has finite downsets, $\operatorname{im}F$ has non-empty intersection with each component of $\operatorname{dn}(q)$. Since $q$ is essential, $\operatorname{dn}(q)$ is disconnected. It follows that $\operatorname{im}\hat{F}=\operatorname{im}F\cap\operatorname{dn}(q)$ is disconnected. Thus, $(F\downarrow q)$ is disconnected by the contrapositive of Proposition˜4.3, a contradiction. We conclude that $E_{Q}\subseteq\operatorname{im}F$, and hence that $I_{Q}\subseteq F(\operatorname{Ob}C)$, establishing (i).
In view of (i), proving (ii) amounts to showing that for each $q\in E_{Q}$ and component $A$ of $\operatorname{dn}(q)$, there is a relation $m<q$ in $\overline{\operatorname{im}F}$ with $m$ a minimum in $A$. Since $Q$ has finite downsets and $M_{Q}\subseteq I_{Q}\subseteq\operatorname{im}F$ by (i), we have $A\cap\operatorname{im}F\neq\emptyset$. Moreover, since $(F\downarrow q)$ is connected, Proposition˜4.3 implies that $\operatorname{im}\hat{F}=\operatorname{im}F\cap\operatorname{dn}[q]$ is connected. Thus, since $q\in\operatorname{im}F$ by (i), there exists at least one relation $a_{0}<q$ in $\operatorname{im}F$ where $a_{0}\in A$. If $a_{0}\in M_{Q}$, we are done. Otherwise, by applying the same argument to any component of $\operatorname{dn}(a_{0})$, we obtain a relation $a_{1}<a_{0}$ in $\operatorname{im}F$ with $a_{1}\in A$. Since $Q$ has finite downsets, by continuing in this way, we eventually obtain a chain of relations
@@
The following example shows that Proposition˜4.4 (ii) is no longer true if we replace $\overline{\operatorname{im}F}$ with $\operatorname{im}F$ in the statement.
@@
Figure 2. The initial functor $F\colon R\to Q$ of Example˜4.5 maps each element of $R$ to the element of $Q$ directly below it. For example, $F(b)=F(c)=x$.
@@
Note that for any category $C$, there is an associated thin category $\operatorname{thin}(C)$ with the same objects such that $\hom(c,c^{\prime})$ is empty in $\operatorname{thin}(C)$ if and only if $\hom(c,c^{\prime})$ is empty in $C$. Moreover, we have a unique functor $\pi\colon C\to\operatorname{thin}(C)$ which is the identity on objects. If $D$ is a thin category (e.g., a poset), then any functor $F\colon C\to D$ induces a functor $\operatorname{thin}(F)\colon\operatorname{thin}(C)\to D$ satisfying $F=\operatorname{thin}(F)\circ\pi$.